OpenStudy (anonymous):

Solve the logarithmic equation. Be sure to reject any value that is not in the domain of the original logarithmic expressions. Give the exact answer. log (x + 29) - log 2 = log (6x + 2)

OpenStudy (agreene):

what base are all of these logs in? or are we just looking for a general answer?

OpenStudy (anonymous):

solve \[\frac{x+29}{2}=6x+2\] and reject any solutions that would make the input negative

OpenStudy (anonymous):

\[x+29=12x+4\] \[33=11x\] \[x=3\]